OpenStudy (anonymous):

A question on thermodynamics of ATP hydrolysis. So, in metabolism of fats, the activation of the fatty acid with CoA, the synthetase enzyme transforms ATP to AMP and pyrophosphate. Then, when we're looking at the energy balance, we assume that two molecules of ATP have been broken down. However, what I found about the Gibbs free energy of ATP hydrolysis to ADP and AMP is respectively -30 and -32 kJ/mol. So, this means that the same amount of energy is released if we break the first or the second anhydride bond in ATP. Why do we assume 2 ATP molecules are broken down in the activation reaction?

OpenStudy (anonymous):

Which is an accurate description of sandy soil?

OpenStudy (aaronq):

I think it's because the regeneration of ATP from AMP takes two individual steps; first to ADP then to ATP, each coupling expends the same amount of energy.

OpenStudy (anonymous):

This is what the textbook says: ATP + H2O → ADP + Pi ΔG = −30 kJ/mol ATP + H2O → AMP + PPi ΔG = −32 kJ/mol So, the -32 kJ/mol corespond to the breakage of one bond, not two

OpenStudy (aaronq):

yeah, but i think the expenditure they talk about is because of the energy needed to regenerate ATP. Cleavage between the \(\alpha\) and \(\beta\) phosphates leaves PPi and AMP, which you would have to add phosphate twice (i.e needs twice as much energy to regenerate ATP). \(AMP+P_i\rightleftharpoons ADP+P_i \rightleftharpoons ATP~~;\Delta G=30+32=62~kJ/mol\) In comparison, cleavage between \(\beta\) and \(\gamma\) phosphates leaves ADP and it needs only one Pi to be converted back (to ATP). \( ADP + P_i\rightleftharpoons ATP~~;\Delta G=30~kJ/mol\)

OpenStudy (anonymous):

Okay, thanks! Makes sense now.

OpenStudy (aaronq):

no problem !